the momentum of particle moving in straight line is given by p=ln t=1 by t kgm by s the time t is greater than 0 at which the net force acting on particle is 0 and its  omentum at that time is

Grade:9

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please upload image of the question and repost as it is not clear what information is given

Regards
Khimraj
4 years ago
p = lnt + 1/t
F = dp/dt = 1/t – 1/t2
F = 0
t2 – t = 0
t = 1
at t=1
p = l1 + 1 = 1
….........................................................